OpenStudy (anonymous):

Find the first four terms of the multiplication pattern given by the formula. 8 • 3^n – 1

OpenStudy (anonymous):

@AloneS

OpenStudy (anonymous):

@Awolflover1

OpenStudy (anonymous):

@jhonyy9

OpenStudy (anonymous):

@hartnn

OpenStudy (yaryar08):

dont you just plug in 1-4 for n?

OpenStudy (anonymous):

so bc. n-1 is the exponent for n=1 what will get ?

OpenStudy (anonymous):

8*3^(1-1) = 8*3^0 = 8*1 = 8 yes ?

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

do you can continue it for n=2 ?

OpenStudy (yaryar08):

plug in 1,2,3,and 4 for the variable n

OpenStudy (anonymous):

oh ok

OpenStudy (yaryar08):

and the results wll be the 1st four terms

OpenStudy (anonymous):

24,72.216,648?

OpenStudy (yaryar08):

well if you plug in 1 you get 8 so it would be 8,24,72,216

OpenStudy (anonymous):

ty

OpenStudy (yaryar08):

np